# The Tantalizing Case of the Quasar J09505128 – I. Presentation of the Data and Detailed Exploration of the Binary Supermassive Black Hole Scenario

### v3

@@ 2.3 Available Light Curves
the Catalina Real-Time Transient Survey (CRTS; drake2009crts), the Near-Earth Object Wide-Field Infrared Survey Explorer (NEOWISE; wright2010wise; mainzer2014wise), and the Zwicky Transient Facility (ZTF; masci2019ztf; bellm2019ztf; graham2019ztf). Figure 2 shows the light curves from these facilities and in their respective filters.
@@ 3 Analysis of Spectra
In Section 3.1, we describe the spectral decomposition and provide central moments and line properties measured from the model broad H lines. Section 3.2 outlines the cross-correlation technique and the resulting velocity shift measurements in both the data and models. In Sections 3.3 and 3.4, we perform simulations to investigate the effects of the observed variability of the broad H profile on the shift measurements. Section 3.5 summarizes the different sets of relative velocity shifts and explains which measurements we consider reliable and adopt for the subsequent radial velocity analysis. Throughout the above analyses we describe how we carry out experiments to determine uncertainties and, in Section 3.6 we summarize those uncertainties.
3.1 Spectral Decomposition and Measurements of the Broad H Line Profiles
We used the PyQSOfit quasar spectrum fitting code (guo2018pyqsofit) to model the spectra and decompose the host galaxy starlight spectrum, quasar continuum, Fe ii lines, broad and narrow He ii 4687 and H 4863 emission lines, and the [O iii] 4959,5007 doublet. The He ii feature is very weak in all the spectra and, although included in the fits, contributes negligibly to the overall models. To perform the decomposition, we restricted the spectral range to 4400–5500 Å and used only the least-squares routine for fitting. Each narrow emission line was modeled with a single Gaussian component, with the centroids and widths of the narrow lines tied together, reflecting their common narrow-line region origin. The broad H line was modeled with four Gaussian components to capture any profile asymmetries and to better fit the less well-defined profiles. The [O iii] wing components were modeled with a Gaussian each and tied to each other in both velocity and width. The width of the Fe ii emission was allowed to vary between 1200–18,000 km s-1. An example decomposition of the 2010 HET spectrum is shown in Figure 3.
@@ 3.1 Spectral Decomposition and Measurements of the Broad H Line Profiles
The broad H profile recovered from a decomposition can be sensitive to the Fe ii fit. To assess this effect, we varied the width of the Fe ii lines in the decomposition process and examined its impact on the modeled broad H profile. We selected a subset of spectra that represent the range of observed broad H profiles of J0950. For each spectrum, we fixed the Fe ii width to be narrower or broader by up to 5000 km s-1 from the best-fit value, ensuring that each fixed width provided a reasonable fit while allowing all other parameters to vary. In each case, we omitted the model Fe ii and all other spectral features from the full model spectrum to isolate the broad H profile. We describe the implications of this Fe ii modeling experiment in Section 3.2.2, where we present the methodology for shift measurements.
@@
3.2 Measuring Broad H Velocity Shifts
The technique used throughout this work to measure shifts of the broad H emission lines of J0950 is the cross-correlation method described in eracleous2012large (see also shen2013constraining). In summary, the method involves first taking two spectra, and linearly scaling one spectrum to match the other in its broad-line flux and continuum. The linear scaling minimizes the difference between the spectra so that the velocity offsets of their broad lines may be compared faithfully. To measure the relative broad H shift, one spectrum is held fixed while the other is progressively shifted in wavelength, and the at each shift is calculated in a window containing only the H profile. The shift value corresponding to the minimum is then adopted. Figure 9 in eracleous2012large demonstrates applications of the cross-correlation method.
@@ 3.2 Measuring Broad H Velocity Shifts
3.2.1 Broad H Velocity Shifts in the Data
We measured velocity shifts of the broad H using two different approaches and compared the results. First, we cross correlated the full spectra in Figure 4, following the same procedure in eracleous2012large, after modeling and subtracting the host galaxy contribution with PyQSOfit (guo2018pyqsofit). In the second approach, we cross correlated the isolated broad H emission lines (Figure 5), obtained by following the spectral decomposition method adopted by other works (shen2013constraining; shen2019sloan; guo2019constraining). To isolate the broad H profiles, we first fit the spectra of J0950 as described in Section 3.1. After fitting, all components except the broad H were subtracted from each spectrum in Figure 4 to get the isolated broad H profiles shown in Figure 5.
@@ 3.2.1 Broad H Velocity Shifts in the Data
As a consistency check, we applied the measured H shifts to four of our spectra that also include broad H (see Figure 1) and verified by visual inspection that those shifts bring the broad H profile into proper alignment. We do not use H as a kinematic tracer because it cannot be observed with H in a single exposure with our primary telescope, the HET, and because H at the redshift of J0950 is contaminated by strong sky lines that ultimately reduce the S/N and undermine our ability to use the narrow lines in its vicinity as a velocity standard. As a second consistency check, we cross-correlated successive spectra and verified that we obtained a shift equivalent to the difference of shifts to the first spectrum. Interestingly, among the observed spectra obtained within two months or less of each other in 2020 and 2022 (see Figure 4), there is a velocity difference of 80 km s-1 between the two spectra in 2022. Such random velocity variations on short timescales are often caused by subtle changes in the broad H line profile, which can be difficult to discern by eye. These velocity fluctuations are manifestations of radial velocity “jitter”, which we discuss further in Sections 3.6 and 5.3.
While the relative velocity shifts from the two methods increase uniformly, the shifts from cross-correlating the isolated H are systematically higher than those from the full spectra at higher velocities. This difference is primarily an artifact caused by the presence of Fe ii emission lines in the full spectra, as discussed in Section 3.2.2.
3.2.2 Broad H Velocity Shifts from Parametric Models of the Data
Using the spectral decomposition method described in Section 3.1, we generated models for the full spectra of J0950. From those, we isolated the broad H models. Models of the full spectra without the contribution from the host galaxy, and isolated broad H lines are shown in Appendix A. Following the methodology in Section 3.2.1, we measured shifts of the broad H using both the full spectra and the isolated broad H emission lines, this time analyzing the models instead of the data.
@@ 3.2.2 Broad H Velocity Shifts from Parametric Models of the Data
Cross-correlating the broad H lines with the two approaches––using models of the full spectra and using models of the broad H profiles only––yielded two sets of relative velocity shifts, shown as a black dotted line and black circles in Figure 6. The two sets of measurements are qualitatively the same, with the velocity shift increasing monotonically in both the full and isolated broad H models. Nevertheless, we performed tests to determine which set of measurements is more reliable. Specifically, we cross correlated each isolated broad H line of the data in Figure 5 with its corresponding model (see Appendix A), which, as expected, yielded zero shift. However, including the Fe ii lines such that the cross-correlation is between an isolated broad H line from the data and its model of broad H plus Fe ii lines, gave non-zero shift values. Therefore, we conclude that the presence of the Fe ii complex introduces a systematic shift in the cross-correlation results. Moreover, the relative velocity shifts derived from the models (black dotted line and black circles in Figure 6) allow us to estimate the measurement uncertainty by carrying out multiple realizations of the noise. For these reasons, we adopt the relative velocity measurements based on the model broad H lines throughout this work. We also perform the same analysis on the relative velocity measurements from the isolated broad H lines of the data, and compare the results (see Section 3.5).
Furthermore, we used the Fe ii width experiment described in Section 3.1 to test how the broadening of the Fe ii template affects the relative velocity shifts obtained by cross correlating the isolated broad H lines. The aim was to assess how uncertainties in the Fe ii modeling would impact shift measurements. We measured the relative velocity shifts under two scenarios described in Section 3.1: one where the Fe ii was modeled as narrower, and another where it was modeled as broader than the best-fit width. The resulting shifts differ by about 10–75 km s-1, with the larger differences corresponding to the broader H profiles seen in the later spectra of J0950. However, these shifts are all consistent within the measurement uncertainties. The effect of varying the width of the Fe ii lines is small because the peak of the broad H, which is well captured in the modeling, contributes more strongly to cross-correlation measurements than the surrounding, less prominent features. Additionally, the Fe ii model fits the region under the broad H used for cross-correlation much better than it fits the strong Fe ii multiplets on either side (see the residuals in Figure 3).
3.3 Simulating the Profile Change of the Broad H
@@
Profile variability is common in normal quasars, but has been known to produce large uncertainties or even invalidate shift measurements (e.g., shen2013constraining; liu2014constraining; runnoe2017large). To test the effects of the observed profile variability of J0950 on our relative shift measurements, we simulated a spectral series based on the model of the broad H line of the 2002 SDSS spectrum. First, we convolved the model with Gaussian kernels of increasing width, and applied velocity shifts to the resulting profiles. Gaussian kernels offer a simple approximation without built-in physical assumptions, and allow us to emulate the evolution of the line profile. As we demonstrate in the next paragraph, using Gaussian kernels does not affect our shift measurements. Following the procedure in Section 3.2.2, we injected synthetic Gaussian noise into each simulated profile.
@@
We show the simulated spectra in Figure 7, with the broadest spectrum having a full width at half maximum (FWHM) comparable to that of the last 2022 spectrum. We then carried out the cross-correlation procedure described in Section 3.2.2 to find the shift of each simulated profile relative to the 2002 broad H profile. The goal of this exercise was to see if cross-correlation can recover the shifts inserted into the simulated profiles. The results are shown in Figure 8, where each shift value corresponds to the median and 68% confidence interval of the distribution resulting from carrying out multiple realizations of noise before cross-correlation. Figure 8 confirms that the input shifts are reliably recovered despite the broadening of the profiles. From this simulation we also obtain an estimate of the typical uncertainty in the velocity shifts arising from the increasing width of the H profile with time: 80 km s-1. It is important to note that these simulated profiles are more round-topped than the actual observed profiles, all of which appear to have cusps (see Figure 5). Therefore, if cross-correlation can recover the shifts of profiles that are more round-topped than those of the data, then shift measurements made for the data are also reliable.
3.4 Simulating Breathing of the Broad-Line Region
@@
We also explored whether the observed variation of J0950’s broad H line can be explained by “breathing” of the BLR (e.g., peterson2011masses; wang2020sloan). Breathing is the effect in which changes in the accretion-powered continuum of a quasar affect the extent of the zone in the BLR that can emit a line efficiently, hence the widths of the resulting broad lines become broader as those lines become fainter. To examine if the broad H behavior of J0950 can be replicated by breathing, and if that has an impact on our shift measurements, we repeated the exercise of Section 3.3 with the difference that we simulated the spectra by applying breathing effects. Specifically, we used a relationship derived under the premise that the observed continuum flux and width of the broad H line are connected to the mass of the BH by (bentz2009radius; peterson2011masses). Since is constant, we can write
@@
This exercise leads us to conclude the following: (a) the decline in integrated flux with increasing width of the broad lines in this simulation (Figure 9) is more drastic than that observed in the broad H line of J0950 (compare Figure 5). Moreover, as noted at the beginning of Section 3.3, the broad H flux of this object exhibits only a 20% stochastic variability––a behavior inconsistent with breathing since the FWHM increases systematically by a factor of over the monitoring period. (b) It is clear from Figure 10 that known shifts from the simulation can be measured despite the effects of breathing, which leads us to infer that symmetric profile variability, in general, has a negligible effect on the shift measurements.
3.5 The Radial Velocity Curve
In Sections 3.2.13.2.2 we obtained four different sets of relative velocity shifts by cross-correlating the broad H lines in: (a) the full spectra of the data, (b) the isolated broad H lines of the data, (c) the full spectra models, and (d) the isolated broad H models. In Section 3.2.2, we established that the presence of Fe ii lines in the full spectra (both data and models), introduces a systematic shift in the cross-correlation results, making shift measurements from the full spectra less reliable.
In view of the above, we proceeded only with the relative shift measurements based on the isolated broad H lines (Figure 5), and their corresponding models (Appendix A). We added to these shift values the peak offset measurement of the broad H of the 2002 spectrum. To measure the peak offset (following a method similar to eracleous2012large), we defined a fixed region around the peak and repeatedly fitted it with a Gaussian, each time varying the wavelength range used for fitting within that region. We took the median peak velocity as the best estimate of the true value and the 68% confidence interval of the resulting distribution of peak values as the measurement uncertainty. It is important to note that this method of locating the peak is sensitive to asymmetries within the broad emission line. Therefore, the Gaussian fitting region must be restricted to be near the peak, excluding significant asymmetries.
The addition of the peak offset from the 2002 spectrum provides the absolute peak offsets of the broad lines, as listed in Table 3. The measurements agree within 3, with average differences around 2. We return to these measurements in Section 5.3 to discuss the implications.

### v4

@@ II.3 Available Light Curves
the Catalina Real-Time Transient Survey (CRTS; Drake et al. 2009), the Near-Earth Object Wide-Field Infrared Survey Explorer (NEOWISE; Wright et al. 2010; Mainzer et al. 2014), and the Zwicky Transient Facility (ZTF; Masci et al. 2019; Bellm et al. 2019; Graham et al. 2019). Figure 2 shows the light curves from these facilities and in their respective filters.
@@ III Analysis of Spectra
In Section III.1, we describe the spectral decomposition and provide central moments and line properties measured from the model broad H lines. Section III.2 outlines the cross-correlation technique and the resulting velocity shift measurements in both the data and models. In Sections III.3 and III.4, we perform simulations to investigate the effects of the observed variability of the broad H profile on the shift measurements. Section III.5 summarizes the different sets of relative velocity shifts and explains which measurements we consider reliable and adopt for the subsequent radial velocity analysis. Throughout the above analyses we describe how we carry out experiments to determine uncertainties and, in Section III.6 we summarize those uncertainties.
III.1 Spectral Decomposition and Measurements of the Broad H Line Profiles
We used the PyQSOfit quasar spectrum fitting code (Guo et al. 2018) to model the spectra and decompose the host galaxy starlight spectrum, quasar continuum, Fe ii lines, broad and narrow He ii 4687 and H 4863 emission lines, and the [O iii] 4959,5007 doublet. The He ii feature is very weak in all the spectra and, although included in the fits, contributes negligibly to the overall models. To perform the decomposition, we restricted the spectral range to 4400–5500 Å and used only the least-squares routine for fitting. Each narrow emission line was modeled with a single Gaussian component, with the centroids and widths of the narrow lines tied together, reflecting their common narrow-line region origin. The broad H line was modeled with four Gaussian components to capture any profile asymmetries and to better fit the less well-defined profiles. The [O iii] wing components were modeled with a Gaussian each and tied to each other in both velocity and width. The width of the Fe ii emission was allowed to vary between 1200–18,000 km s-1. An example decomposition of the 2010 HET spectrum is shown in Figure 3.
@@ III.1 Spectral Decomposition and Measurements of the Broad H Line Profiles
The broad H profile recovered from a decomposition can be sensitive to the Fe ii fit. To assess this effect, we varied the width of the Fe ii lines in the decomposition process and examined its impact on the modeled broad H profile. We selected a subset of spectra that represent the range of observed broad H profiles of J0950. For each spectrum, we fixed the Fe ii width to be narrower or broader by up to 5000 km s-1 from the best-fit value, ensuring that each fixed width provided a reasonable fit while allowing all other parameters to vary. In each case, we omitted the model Fe ii and all other spectral features from the full model spectrum to isolate the broad H profile. We describe the implications of this Fe ii modeling experiment in Section III.2.2, where we present the methodology for shift measurements.
@@
III.2 Measuring Broad H Velocity Shifts
The technique used throughout this work to measure shifts of the broad H emission lines of J0950 is the cross-correlation method described in Eracleous et al. (2012) (see also Shen et al., 2013). In summary, the method involves first taking two spectra, and linearly scaling one spectrum to match the other in its broad-line flux and continuum. The linear scaling minimizes the difference between the spectra so that the velocity offsets of their broad lines may be compared faithfully. To measure the relative broad H shift, one spectrum is held fixed while the other is progressively shifted in wavelength, and the at each shift is calculated in a window containing only the H profile. The shift value corresponding to the minimum is then adopted. Figure 9 in Eracleous et al. (2012) demonstrates applications of the cross-correlation method.
@@ III.2 Measuring Broad H Velocity Shifts
III.2.1 Broad H Velocity Shifts in the Data
We measured velocity shifts of the broad H using two different approaches and compared the results. First, we cross correlated the full spectra in Figure 4, following the same procedure in Eracleous et al. (2012), after modeling and subtracting the host galaxy contribution with PyQSOfit (Guo et al. 2018). In the second approach, we cross correlated the isolated broad H emission lines (Figure 5), obtained by following the spectral decomposition method adopted by other works (Shen et al. 2013, 2019; Guo et al. 2019). To isolate the broad H profiles, we first fit the spectra of J0950 as described in Section III.1. After fitting, all components except the broad H were subtracted from each spectrum in Figure 4 to get the isolated broad H profiles shown in Figure 5.
@@ III.2.1 Broad H Velocity Shifts in the Data
As a consistency check, we applied the measured H shifts to four of our spectra that also include broad H (see Figure 1) and verified by visual inspection that those shifts bring the broad H profile into proper alignment. We do not use H as a kinematic tracer because it cannot be observed with H in a single exposure with our primary telescope, the HET, and because H at the redshift of J0950 is contaminated by strong sky lines that ultimately reduce the S/N and undermine our ability to use the narrow lines in its vicinity as a velocity standard. As a second consistency check, we cross-correlated successive spectra and verified that we obtained a shift equivalent to the difference of shifts to the first spectrum. Interestingly, among the observed spectra obtained within two months or less of each other in 2020 and 2022 (see Figure 4), there is a velocity difference of 80 km s-1 between the two spectra in 2022. Such random velocity variations on short timescales are often caused by subtle changes in the broad H line profile, which can be difficult to discern by eye. These velocity fluctuations are manifestations of radial velocity “jitter”, which we discuss further in Sections III.6 and V.3.
While the relative velocity shifts from the two methods increase uniformly, the shifts from cross-correlating the isolated H are systematically higher than those from the full spectra at higher velocities. This difference is primarily an artifact caused by the presence of Fe ii emission lines in the full spectra, as discussed in Section III.2.2.
III.2.2 Broad H Velocity Shifts from Parametric Models of the Data
Using the spectral decomposition method described in Section III.1, we generated models for the full spectra of J0950. From those, we isolated the broad H models. Models of the full spectra without the contribution from the host galaxy, and isolated broad H lines are shown in Appendix A. Following the methodology in Section III.2.1, we measured shifts of the broad H using both the full spectra and the isolated broad H emission lines, this time analyzing the models instead of the data.
@@ III.2.2 Broad H Velocity Shifts from Parametric Models of the Data
Cross-correlating the broad H lines with the two approaches––using models of the full spectra and using models of the broad H profiles only––yielded two sets of relative velocity shifts, shown as a black dotted line and black circles in Figure 6. The two sets of measurements are qualitatively the same, with the velocity shift increasing monotonically in both the full and isolated broad H models. Nevertheless, we performed tests to determine which set of measurements is more reliable. Specifically, we cross correlated each isolated broad H line of the data in Figure 5 with its corresponding model (see Appendix A), which, as expected, yielded zero shift. However, including the Fe ii lines such that the cross-correlation is between an isolated broad H line from the data and its model of broad H plus Fe ii lines, gave non-zero shift values. Therefore, we conclude that the presence of the Fe ii complex introduces a systematic shift in the cross-correlation results. Moreover, the relative velocity shifts derived from the models (black dotted line and black circles in Figure 6) allow us to estimate the measurement uncertainty by carrying out multiple realizations of the noise. For these reasons, we adopt the relative velocity measurements based on the model broad H lines throughout this work. We also perform the same analysis on the relative velocity measurements from the isolated broad H lines of the data, and compare the results (see Section III.5).
Furthermore, we used the Fe ii width experiment described in Section III.1 to test how the broadening of the Fe ii template affects the relative velocity shifts obtained by cross correlating the isolated broad H lines. The aim was to assess how uncertainties in the Fe ii modeling would impact shift measurements. We measured the relative velocity shifts under two scenarios described in Section III.1: one where the Fe ii was modeled as narrower, and another where it was modeled as broader than the best-fit width. The resulting shifts differ by about 10–75 km s-1, with the larger differences corresponding to the broader H profiles seen in the later spectra of J0950. However, these shifts are all consistent within the measurement uncertainties. The effect of varying the width of the Fe ii lines is small because the peak of the broad H, which is well captured in the modeling, contributes more strongly to cross-correlation measurements than the surrounding, less prominent features. Additionally, the Fe ii model fits the region under the broad H used for cross-correlation much better than it fits the strong Fe ii multiplets on either side (see the residuals in Figure 3).
III.3 Simulating the Profile Change of the Broad H
@@
Profile variability is common in normal quasars, but has been known to produce large uncertainties or even invalidate shift measurements (e.g., Shen et al. 2013; Liu et al. 2014; Runnoe et al. 2017). To test the effects of the observed profile variability of J0950 on our relative shift measurements, we simulated a spectral series based on the model of the broad H line of the 2002 SDSS spectrum. First, we convolved the model with Gaussian kernels of increasing width, and applied velocity shifts to the resulting profiles. Gaussian kernels offer a simple approximation without built-in physical assumptions, and allow us to emulate the evolution of the line profile. As we demonstrate in the next paragraph, using Gaussian kernels does not affect our shift measurements. Following the procedure in Section III.2.2, we injected synthetic Gaussian noise into each simulated profile.
@@
We show the simulated spectra in Figure 7, with the broadest spectrum having a full width at half maximum (FWHM) comparable to that of the last 2022 spectrum. We then carried out the cross-correlation procedure described in Section III.2.2 to find the shift of each simulated profile relative to the 2002 broad H profile. The goal of this exercise was to see if cross-correlation can recover the shifts inserted into the simulated profiles. The results are shown in Figure 8, where each shift value corresponds to the median and 68% confidence interval of the distribution resulting from carrying out multiple realizations of noise before cross-correlation. Figure 8 confirms that the input shifts are reliably recovered despite the broadening of the profiles. From this simulation we also obtain an estimate of the typical uncertainty in the velocity shifts arising from the increasing width of the H profile with time: 80 km s-1. It is important to note that these simulated profiles are more round-topped than the actual observed profiles, all of which appear to have cusps (see Figure 5). Therefore, if cross-correlation can recover the shifts of profiles that are more round-topped than those of the data, then shift measurements made for the data are also reliable.
III.4 Simulating Breathing of the Broad-Line Region
@@
We also explored whether the observed variation of J0950’s broad H line can be explained by “breathing” of the BLR (e.g., Peterson 2011; Wang et al. 2020). Breathing is the effect in which changes in the accretion-powered continuum of a quasar affect the extent of the zone in the BLR that can emit a line efficiently, hence the widths of the resulting broad lines become broader as those lines become fainter. To examine if the broad H behavior of J0950 can be replicated by breathing, and if that has an impact on our shift measurements, we repeated the exercise of Section III.3 with the difference that we simulated the spectra by applying breathing effects. Specifically, we used a relationship derived under the premise that the observed continuum flux and width of the broad H line are connected to the mass of the BH by (Bentz et al. 2009; Peterson 2011). Since is constant, we can write
@@
This exercise leads us to conclude the following: (a) the decline in integrated flux with increasing width of the broad lines in this simulation (Figure 9) is more drastic than that observed in the broad H line of J0950 (compare Figure 5). Moreover, as noted at the beginning of Section III.3, the broad H flux of this object exhibits only a 20% stochastic variability––a behavior inconsistent with breathing since the FWHM increases systematically by a factor of over the monitoring period. (b) It is clear from Figure 10 that known shifts from the simulation can be measured despite the effects of breathing, which leads us to infer that symmetric profile variability, in general, has a negligible effect on the shift measurements.
III.5 The Radial Velocity Curve
In Sections III.2.1III.2.2 we obtained four different sets of relative velocity shifts by cross-correlating the broad H lines in: (a) the full spectra of the data, (b) the isolated broad H lines of the data, (c) the full spectra models, and (d) the isolated broad H models. In Section III.2.2, we established that the presence of Fe ii lines in the full spectra (both data and models), introduces a systematic shift in the cross-correlation results, making shift measurements from the full spectra less reliable.
In view of the above, we proceeded only with the relative shift measurements based on the isolated broad H lines (Figure 5), and their corresponding models (Appendix A). We added to these shift values the peak offset measurement of the broad H of the 2002 spectrum. To measure the peak offset (following a method similar to Eracleous et al. 2012), we defined a fixed region around the peak and repeatedly fitted it with a Gaussian, each time varying the wavelength range used for fitting within that region. We took the median peak velocity as the best estimate of the true value and the 68% confidence interval of the resulting distribution of peak values as the measurement uncertainty. It is important to note that this method of locating the peak is sensitive to asymmetries within the broad emission line. Therefore, the Gaussian fitting region must be restricted to be near the peak, excluding significant asymmetries.
The addition of the peak offset from the 2002 spectrum provides the absolute peak offsets of the broad lines, as listed in Table 3. The measurements agree within 3, with average differences around 2. We return to these measurements in Section V.3 to discuss the implications.